OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1) 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6 2) 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

i did 2 i just need one now

OpenStudy (amistre64):

if its true for all positive ints, then test it for n=1

OpenStudy (amistre64):

if its true for n=1, then assume its true for all n=k

OpenStudy (amistre64):

im not too sure why "assuming" something to be true actually makes it true, but for some reason thats how this goes

OpenStudy (anonymous):

so 4(1)(4(1)+2)/2 = 24/2= 12 not one

OpenStudy (anonymous):

so it's not true?

OpenStudy (amistre64):

\[P(1)=4(1)( 4(1) + 2) = \frac{4(4(1)+1)(8(1)+7)}{6}\] \[P(1)=4\cdot 6 = \frac{10(5)}{1}\] is your setup correct?

OpenStudy (amistre64):

well, if its setup correctly, and 1 IS a positive integer .... then its not true for n=1

OpenStudy (anonymous):

oh duh, yes and so does that mean you can't do the problem?

OpenStudy (amistre64):

it means that it false since it states that it is true for ALL positive integers

OpenStudy (anonymous):

Okay thank you very much :)

OpenStudy (amistre64):

youre welcome